OpenStudy (anonymous):

y= 3/4x-2 how do I get y-intercept and x-intercept??

OpenStudy (saifoo.khan):

When x intercpts, y = 0 When y intercpts x = 0

OpenStudy (anonymous):

who to get 0???

OpenStudy (anonymous):

i mean how

OpenStudy (saifoo.khan):

Like then you want to get y intercept, put x the zero. y = 3x/4 -2 y = 3(0)/4 - 2 y = 0 -2 y = 2 So, (0,2)

OpenStudy (anonymous):

why r u multiply 0 nd where did u get 0 from??

OpenStudy (saifoo.khan):

i made x zero. this is b/c when a line cuts the y axis, the x is always zero.

OpenStudy (saifoo.khan):